Find the center, radius, intercepts and graph: x^2+y^2+10x-4y+4=0
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Find the center, radius, intercepts and graph: x^2+y^2+10x-4y+4=0
----------
x^2+y^2+10x-4y = -4
(x^2+10x+0) + (y^2-4y+0) = -4
Complete the squares for x & y

Answer by ikleyn(52846)

.
If you don't know how to complete the squares and to make the required transformations, see the lessons
    - Standard equation of a circle
    - General equation of a circle
    - Transform general equation of a circle to the standard form by completing the squares
    - Identify elements of a circle given by its general equation
in this site.

You will find there the theory, examples and complete and detailed explanations and illustrations.


Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic
"Conic sections: Ellipses. Definition, major elements and properties. Solved problems".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.